OpenStudy (anonymous):

i really need help with this my paper is due tomorrow/ one edge s of a cube is 3 cm long. substituting 3 in the volume formula, v=s3, gives the volume of the cube:v=3.the square having a side s that is 3 cm long. substituting 3 in the area formula,A=s2, gives the area of the square :a=3to the 2nd power. i need help dividing this i need to divide the volume of the cube by the area of the square

OpenStudy (bahrom7893):

What?

OpenStudy (anonymous):

hold on

OpenStudy (bahrom7893):

One edge of a cube is 3cm long, therefore: V(cube) = 3^3 = 27 One side of a square is 3 cm long, therefore: A(square) = 3^2 = 9 V(cube)/A(square) = 27/9 = 3

OpenStudy (bahrom7893):

so is that it?

OpenStudy (anonymous):

i think so

OpenStudy (bahrom7893):

k im good, lol i always come here to ask my question, nobody can solve it cuz my math is too advanced, so i end up helpin others lol

OpenStudy (anonymous):

kool

OpenStudy (anonymous):

i have a lot of problems in math

OpenStudy (anonymous):

k12/k11,k=0

OpenStudy (bahrom7893):

what?

OpenStudy (anonymous):

?

OpenStudy (bahrom7893):

what do u mean with that question?

OpenStudy (anonymous):

never mind i figured it out

OpenStudy (anonymous):

im stuck on equations x+b=c and i have some equations and story problems that go with it im in algebra 1 9th grade

OpenStudy (bahrom7893):

k if x+b = c then x = c - b

OpenStudy (anonymous):

yeah thats just the formula /x+(-9)=3 this is the problem

OpenStudy (bahrom7893):

x = 3 - (-9) x = 3+9 = 12

OpenStudy (anonymous):

i have 11 more problems like this

OpenStudy (anonymous):

\[d+(-3)=0\]

OpenStudy (bahrom7893):

d = 3

OpenStudy (bahrom7893):

and thats basic u should do it on ur own

OpenStudy (anonymous):

ok im starting to get it more thank you